The hydrogen gas needed to make ammonia, hydrogen chloride gas, and methanol can be obtained from small hydrocarbons such as methane or propane through the steam-reforming process, conducted at 41 atm and 760-980 °C. If 2.7 × 107 L of C3H8(g) at 810 °C and 8.0 atm react in the first step of the process, shown below, what is the maximum volume in liters of CO(g) at STP that can form? C3H8 + 3H2O → 3CO + 7H2 3CO + 3H2O → 3CO2 + 3H2
Answer #1

PV = nRT

n = PV/(RT)

n = (8)(2.7*10^7)/((0.082)(810+273)) = 2432268.08999

then

1 mol of fuel = 3 mol of CO

2432268.08999 mol = 2432268.08999*3 = 7296804.26997 mol of CO

at STP

1 mol = 22.4 L

then

7296804.26997 mol = 22.4*7296804.26997 = 163448415.647 L
Answer #2
The hydrogen gas used to make ammonia can be made from small hydrocarbons such as methane in the so-called steam-reforming process, run at high temperature and pressure. a. If 3.2 atm of methane at 21 °C are introduced into a container, to what temperature must the gas be heated to increase the pressure to 12 atm? b. If 4.0 × 103 L of methane gas at 21 °C is heated and allowed to expand at a constant pressure, what will the volume become in m3 when the temperature reaches 815 °C? c. What volume of methane gas at 21 °C and 1.1 atm must be compressed to yield 1.5 × 104 L of methane gas at 12 atm and 815 °C?
source: Fiverr
answered by: anonymous